OpenStudy (anonymous):

A 30 kg kid on rollerskates is moving 1 m/s. His brother pushes him with a force of 5N. What will be his kinetic energy after his brother pushes him 4.0m. (I'm not sure what equation/formula to use)

OpenStudy (anonymous):

What Are We Looking For (WALF): The answer to: \[K=\frac{1}{2}mv^2\] To do this, we need the values of mass (which the question has given us) and velocity (which we will obtain)  So we have: \[m=30kg ,and, v_0=1\frac{m}{s}\] His brother gives him a push: \[F=ma\] So we can say: \[(5N)=(30kg)(a)\] We calculate acceleration as: \[a=\frac{1}{6}\] We now have the square of the velocity and mass: can you give the answer? Using \[v^2=2a(\Delta x)+v^2_0\] We can find out the speed at at 4 meters. So \[v^2=2(\frac{1}{6})(4-0)+1^2\] \[v^2=\frac{7}{3}\]